OpenStudy (anonymous):

find the values of a h and k that make the equation 1/2x^2-3x+5=a(x-h)+k

OpenStudy (anonymous):

ok here we go. best way to do this is compute \[\frac{-b}{2a}\]

OpenStudy (anonymous):

it is going to be \[\frac{1}{2}(x-h)^2+k\]

OpenStudy (anonymous):

im confused is it that simple?

OpenStudy (anonymous):

\[h=\frac{b}{2a}=\frac{-3}{2\times \frac{1}{2}}=-3\]

OpenStudy (anonymous):

so you will have \[\frac{1}{2}(x-3)^2+k\]

OpenStudy (anonymous):

mmhmm

OpenStudy (anonymous):

now to find k, you see that if you replace x by 3 is (x-3) you will get 0+k=k so replace x by 3 in the original equation to find k

OpenStudy (anonymous):

yes it is that simple. usually made to seem harder than it is

OpenStudy (anonymous):

\[h=\frac{b}{2a}\] and to find k replace x by \[-\frac{b}{2a}\] in the original expression

OpenStudy (anonymous):

this is how you find the vertex

OpenStudy (anonymous):

you have \[y=\frac{1}{2}x^2-3x+5\] \[y=\frac{1}{2}3^3-3\times 3+5\] \[y=\frac{9}{2}-9+5\] \[y=-\frac{9}{2}+5\] \[y=\frac{1}{2}\]

OpenStudy (anonymous):

so your answer is \[\frac{1}{2}(x-3)^2+\frac{1}{2}\]

OpenStudy (anonymous):

this is a lot easier than completing the square and trying to figure out what you added and subtracted. it is a straightforward computation

OpenStudy (anonymous):

easy yes? i mean relatively easy